### v4

@@ Abstract
Determining when people are struggling from video enables a finer-grained understanding of actions and opens opportunities for building intelligent support visual interfaces. In this paper, we present a new dataset with three assembly activities and corresponding performance baselines for the determination of struggle from video. Three real-world problem-solving activities including assembling plumbing pipes (Pipes-Struggle), pitching camping tents (Tent-Struggle) and solving the Tower of Hanoi puzzle (Tower-Struggle) are introduced. Video segments were scored w.r.t. the level of struggle as perceived by annotators using a forced choice 4-point scale. Each video segment was annotated by a single expert annotator in addition to crowd-sourced annotations. The dataset is the first struggle annotation dataset and contains 5.1 hours of video and 725,100 frames from 73 participants in total. We evaluate three decision-making tasks: struggle classification, struggle level regression, and struggle label distribution learning. We provide baseline results for each of the tasks utilising several mainstream deep neural networks, along with an ablation study and visualisation of results. Our work is motivated toward assistive systems that analyze struggle, support users during manual activities and encourage learning, as well as other video understanding competencies.
@@
Computer Vision Dataset, Egocentric Video, Struggle Determination
@@ 1 Introduction
Understanding when someone is struggling is an important cognitive competence and a key component of skill acquisition [38]. This capability could enable the development of systems that understand how effectively people perform activities and assist with achieving goals or supporting learning outcomes. The ability to detect struggle and subsequently offer help, is a core developmental ability and exists already in pre-verbal children and even other primates [59].
@@
Motivated by the larger goal of finer-grained visual understanding and opportunities on eye-wear computing to provide real-time assistance, we collected a dataset for struggle determination using indoor and outdoor real-world activities, including three domains: assembling plumbing pipes; pitching a camping tent; and solving the ‘Tower of Hanoi’ puzzle. These three domains were chosen specifically to demonstrate the viability of struggle detection, due to their task structures and assembly requirements. The Tower of Hanoi puzzle presented a strict rule-based framework requiring participants to solve a spatial puzzle. Pipe plumbing uses rigid objects and printed instructions for participants to follow, showcasing their ability to perform an unknown task with guidance. The tent assembly task introduced a free-form assembly process with deformable objects, demanding greater improvisation and manual skill. Compared to other general real daily-living tasks (eg cooking or crafts), our chosen tasks are examples of real-world problem-solving and have well-defined goals (completion of the structures). Our tasks allow participants to demonstrate a variety of complex visuomotor behaviours with multiple, albeit constrained, opportunities for struggle.
@@
Our struggle-determination approach is based on two main assumptions. First, we believe there is sufficient information for a large number of activities to determine if a person is struggling from video, based on low-level motion cues of hand-object interaction and without requiring specific task knowledge. This is supported by behavioural evidence e.g. [59, 60] in people. Second, the same deep learning models and train/test strategies can be used for different tasks as they share common visual cues that help to determine struggle, such as the smoothness of hand motion.
@@
Figure 1 shows sample video frames from our data. We used two levels for annotation: MTurk workers and a domain expert. Annotations used a forced choice, four-level struggle scale so that we can determine if a person is struggling and to what degree. We evaluated several video understanding methods trained on our dataset and provided a variety of baseline results across our proposed struggle determination activities and machine learning tasks.
Our primary contributions are: 1) Surfacing the problem of struggle determination for finer-grained assembly understanding in egocentric videos. 2) Introducing a public dataset for struggle determination, which is annotated via an expert annotator and crowdsourcing. 3) Presenting three benchmark tasks for struggle determination: classification, regression, and label-distributed learning, all using weakly supervised learning with only video-level labels. 4) We provide initial benchmark results of video understanding models on struggle determination thus providing baseline results and demonstrating the essential role of motion information.
@@ 2 Related Work
Numerous deep architectures have been developed for AR. Initially, 2D-ConvNets were created to distinguish video actions from 2D spatial features in video frames, typically combined with temporal average pooling and/or optical flow to improve motion understanding. Subsequently, 3D-ConvNets [55, 4, 15, 61, 56, 57, 16] gained popularity due to their ability to model motion in the 3D convolution operation, as well as being able to train using large-scale datasets, such as Kinetics [25] and the proposal of expanding 2D convolutions into 3D [4].
Recently, Transformer-based architectures such as Vision Transformer (ViT) [8], Swin Transformer [33], MLP-Mixer [54], and gMLP [30] have been used increasingly. For AR, a hierarchical architecture has been developed to extract spatial-temporal features efficiently, such as ViViT [1], TimeSformer [3], MViT [14, 28], and Video Transformer Networks (VTN) [37].
@@
Action Quality Assessment. Action quality assessment (AQA) is a regression task predicting action quality scores. Most datasets, such as AQA-7 [41], Olympic Scoring Dataset [42, 44], JIGSAW [18], and MTL-AQA [43], contain sports scenes, where annotations represent the scores given by expert human judges. The typical architecture for AQA consists of 3D convolutional neural networks (C3D) [55] to extract spatial-temporal features, followed by average pooling (C3D-AVG), LSTM [22] (C3D-LSTM), or attention modules to further aggregate information from various video segments and generate regression or multitask predictions [43]. Some methods tackle this task using Label Distribution Learning [51] in which the distribution of assessment scores is predicted. Methods often predict the scores directly using MLPs [65], or the mean/std. dev of a Gaussian Distribution based on Variational Auto-Encoders (VAE) [51, 68].
@@
Skill Determination. Similar to AQA, Skill Determination [40, 39, 9, 10, 21, 29] aims to predict the skill level of the action being performed but does so via ranking. [40, 39] mainly focus on ranking surgery skills using Siamese LSTM-based networks by aggregating kinematic data. [9] investigated ranking skills for general daily activities and introduced the EPIC-Skill2018 dataset. [29] proposes an RNN-based spatial attention model and significantly improves the results in [9]. [10] employs a combination of contrastive learning and attention to discriminate between high- and low-skill actions that require fine-grained discrimination skills. Most recently, [21] proposes a TikTok dance video dataset for ranking dancing performances.
@@
Video Anomaly Action Detection. Video Anomaly Detection aims to detect unintentional [11, 64] or anomalous actions [69, 49, 53, 70]. Whilst fine-grained discrimination may share some similarities to Struggle Detection we differ in two major aspects: persons struggling may not contain anomalies and Anomaly Detection is constructed as a temporal localisation task.
@@
We find that fine-grained video understanding has been explored in a variety of different datasets and tasks (see Table 6 in the Appendix). However, albeit crucial for fine-grained activity and skill understanding, determining the level of struggle of the participant has not been attempted. Previous research efforts motivate us to create three tasks for our dataset: struggle classification, struggle level regression, and struggle label distribution learning. As labels for these tasks do not exist, we annotate struggle for one existing dataset [23, 48] and collect and annotate two new activities. Our three activities were chosen to require complex visuomotor coordination and problem-solving while remaining goal-driven. This promotes similar yet varied occurrences of struggle behaviours and provides enough diversity within and across activities.
@@ 3 Struggle Determination Dataset
We initially recorded the behaviour of the participants in three scenarios that require complex visuomotor coordination and problem-solving and where participants may occasionally make errors or otherwise struggle. These scenarios were intended to be similar to what one might encounter in daily life (both indoors and outdoors) and where one might want to receive instructional advice. Finding suitable scenarios for struggle detection presents a balancing act. The scenarios shouldn’t be too easy, as participants might not exhibit struggle behaviours, but they also shouldn’t be overly complex or open-ended to prevent unclear goals and excessive task completion times (we favour activities contained within a few minutes). One alternative option would have been to leverage existing visuomotor datasets, such as EPIC-Kitchen [7], by selecting and annotating videos related to struggle. We have investigated this option and generated Table 6 (see in appendix) summarizing potential datasets. Unfortunately, we encountered two key limitations on most previous datasets: (1) Misaligned annotations: None of the existing datasets offered annotations that specifically targeted or purposefully quantified struggle; (2) Scarcity of repeated goal-directed tasks: existing datasets mostly lack repeated instances of goal-directed activities (e.g., not every participant is cracking eggs or flipping pancakes). These limitations led us to consider collecting new data rather than just relying on the existing ones. An exception is [23, 48], where several subjects repeat the same activity for a complex visuomotor assembly of a camping tent. We annotate this data in our work. Finally, we want real scenarios that are ordinary enough so that naive observers can estimate if a participant is struggling without prior knowledge. Across the three activities, we recorded first-person videos and later supplemented the videos with struggle-determination annotations from human observers.
@@
The scenarios included: assembling plastic plumbing pipes by following one of two diagrams (easy or difficult) while seated indoors; solving a version of the Tower of Hanoi puzzle while seated indoors; and finally assembling a camping tent moving freely outdoors. We respectively name the scenarios as Pipes, Tower, and Tent and the corresponding activities/subsets for struggle determination as Pipes-Struggle, Tower-Struggle and Tent-Struggle. Each scenario includes fine-grained actions that people perform routinely (e.g. arranging, picking up and placing objects, attaching them, etc.). Equipment needed, difficulty and goals vary between scenarios, as does participant performance, with some struggling more or less in completing a scenario.
@@
Slightly non-struggling The actions are generally executed smoothly and with confidence, with only occasional minor errors or corrections that indicate some level of challenge or difficulty.
Slightly struggling There are some signs of challenge or difficulty including occasional errors, pauses, dropping off items, or wobbly hands that indicate the participants may not be very experienced in the task at hand but are still making an effort to complete it.
@@ 3.1.1 Data Collection
For all three scenarios, university undergraduate and graduate student populations were recruited. All participants gave informed consent and approval for data sharing. We recruited 24 individuals to perform the plumbing task (80% female, mean age 22.6 years), and 20 participants from this group also performed the Tower of Hanoi task. For Struggle-Tent, as stated in [23], 29 individuals (71% female, mean age 23.3 years) performed the tent-pitching scenario. In all activities, participants wore a head-mounted GoPro Hero 5 (1080p at 60Hz).
@@
In the plumbing task, participants assembled home plumbing pipes following a diagram of a double sink layout (see supplementary material). Participants were evenly split between those with an easier and more difficult diagram to follow. To simplify dexterity requirements, the activity was reduced to assembling the layout of the pipes on a table rather than in a real sink.
Videos were captured at $1920\times 1080$ resolution, at 30Hz for Tent-Struggle, and 60Hz for Pipes-Struggle and Tower-Struggle. Struggle determination annotation was accomplished by MTurkers rating 10s video clips from the tasks. We segment videos into 10s blocks. Given the struggle’s inherent ambiguity, we discovered during the early annotation process that it can be difficult for annotators to locate the beginning and end of the struggle in untrimmed video recordings. Therefore we chose to use a weakly labelling approach instead. Inspired by Experimental Psychology research (e.g. [59], [60]) that shows pre-verbal children and even chimpanzees can determine struggle actions within 10s, our clips were created by uniformly splitting the videos into 10-second segments from beginning to end (e.g. 0-9s, 10-19s, etc.). We also resized videos to $456\times 256$ at 30Hz. This ensured segments had a sufficient resolution for human observers while maintaining a resolution suitable for deep learning models. Our annotation process thus focuses on evaluating struggle within isolated 10-second clips. This means that we assess and label the presence of struggle solely within each clip, without considering the broader context of the full video recording. Using the EPIC-Tent action label annotations [23], we chose to focus on video segments from five actions (assemble support, insert stake, insert support, insert support tab, and place guyline) where our expert annotator judged they had the most obvious visual cues of struggle in a series of pilot studies.

### v5

@@ Abstract
Determining when people are struggling allows for a finer-grained understanding of actions that complements conventional action classification and error detection. Struggle detection, as defined in this paper, is a distinct and important task that can be identified without explicit step or activity knowledge. We introduce the first struggle dataset with three real-world problem-solving activities that are labelled by both expert and crowd-source annotators. Video segments were scored w.r.t. their level of struggle using a forced choice 4-point scale. This dataset contains 5.1 hours of video from 73 participants. We conducted a series of experiments to identify the most suitable modelling approaches for struggle determination. Additionally, we compared various deep learning models, establishing baseline results for struggle classification, struggle regression, and struggle label distribution learning. Our results indicate that struggle detection in video can achieve up to $88.24\%$ accuracy in binary classification, while detecting the level of struggle in a four-way classification setting performs lower, with an overall accuracy of $52.45\%$. Our work is motivated toward a more comprehensive understanding of action in video and potentially the improvement of assistive systems that analyse struggle and can better support users during manual activities.
@@
Struggle Detection, Struggle Dataset, Egocentric Problem-Solving Videos, Deep Learning Models
@@ 1 Introduction
Figure 2: Struggle vs Errors/Mistakes. The figure showcases the differences between error detection and struggle determination using video examples from our dataset: (1) participant showing no sign of struggle, but there was an error in the completed activity; (2) participant showing signs of struggle but completed the activity successfully without any errors; and (3) participant showing signs of struggle with an error in the completed activity.
The ability to identify when someone is struggling is an important cognitive competence and a key component of skill acquisition [44]. Observers can identify instances of struggle even when lacking specific task knowledge (e.g. a non-rock climber can determine when expert rock climbers are struggling). The ability to detect struggle and subsequently offer help has also been demonstrated in pre-verbal children and even other primates [70]. These findings suggest that struggle determination is a rather important and fundamental ability of visual action understanding which we define as a task within this paper—see examples within Figure 1. However, most research to date in skill assessment has focused on error detection [20, 69, 54, 23, 17], ranking of skill [9, 10] or task scoring [48, 50]. Struggle, however, is distinct from error detection and these other tasks. It is possible to complete activities without step mistakes while struggling or to confidently complete an activity full of step mistakes without showing signs of struggle. Figure 2 presents examples from our dataset illustrating struggle-error relations. Struggle and error detection, though related, are distinct markers of action assessment. Importantly, the struggle is detectable without prior task knowledge, can be used to anticipate errors, can serve as an online scoring predictor before tasks are completed, and can assist an assistive system in identifying common steps that are difficult to perform. Therefore, we argue that visual understanding systems should incorporate both error and struggle detection. Struggle, a subjective behavioural marker, is challenging to define precisely. Nevertheless, it is easily detectable by non-experts, allowing for the construction of relevant datasets. In this work, we define struggle through the observations of annotators, where we observe that instances of struggle are identified during periods of motor hesitation (e.g. stopping or showing placement indecision), getting stuck or not knowing how to proceed, actions taking longer than expected; having non-smooth hand or body motions; frequent pauses and repeated attempts; or showing signs of frustration (e.g. through hand and or head movements).
@@
Motivated by the importance of struggle determination within the larger goal of finer-grained visual understanding and the potential opportunities for eye-wear computing to provide real-time support assistance, we collected a dataset for struggle determination using indoor and outdoor real-world activities, including three domains: assembling plumbing pipes; pitching a camping tent; and solving the ‘Tower of Hanoi’ puzzle. These three domains were chosen specifically to demonstrate the viability of struggle detection due to their task structures and assembly requirements. The Tower of Hanoi puzzle presented a strict rule-based framework requiring participants to solve a spatial puzzle. Pipe plumbing uses rigid objects and printed instructions for participants to follow, showcasing their ability to perform an unknown task with guidance. The tent assembly task introduced a free-form assembly process with deformable objects, demanding greater improvisation and manual skill. Our chosen tasks are examples of real-world problem-solving that require structured, sequential assembly with a predefined final state. These tasks involve following specific rules and constraints to achieve a well-defined outcome, such as correctly connecting components or completing a structured puzzle. Our tasks allow participants to demonstrate various complex visuomotor behaviours with multiple, albeit constrained, opportunities for struggle.
@@
Our struggle-determination approach is based on two main assumptions. First, we believe there is sufficient information for many activities to determine if a person is struggling from video, based on low-level motion cues of hand-object interaction and without requiring specific task knowledge. This is supported by behavioural evidence, e.g. [70, 71] in people. Second, the same deep learning models and related training and testing strategies can be used for different tasks as they share common visual cues that help to determine struggle, such as the smoothness of hand motion.
@@
Figure 1 shows sample video frames from our data. We used two levels for annotation: MTurk workers and a domain expert. Annotations used a forced choice, four-level struggle scale so that we can determine if a person is struggling and to what degree. We evaluated several video understanding methods trained on our dataset and provided various baseline results across our proposed struggle determination activities and machine learning tasks.
In summary, our contributions are three-fold:
  1. 1.
    We define the problem of struggle determination as a fine-grained task from egocentric assembly videos and introduce a public dataset for struggle determination. This includes new data subsets for activities Pipes-Struggle and Tower-Struggle, as well as struggle annotation for activity Tent-Struggle from EPIC-Tent [26]. All data is annotated using both an expert annotator and crowd-sourcing.
  2. 2.
    We conducted experiments to identify the most suitable modelling approach for struggle determination and to compare the performance of different deep models to determine the most efficient one. These results serve as baseline benchmarks for struggle determination, not only as a classification task but also through regression and label distribution learning as alternative approaches.
  3. 3.
    We demonstrate the essential role of motion information in struggle determination and present preliminary results on the generalization across the three activities.
@@ 2 Related Work
Numerous deep architectures have been developed for action recognition (AR). Initially, 2D-ConvNets were designed to distinguish video actions from 2D spatial features in video frames. These models typically used temporal average pooling and/or optical flow to enhance motion understanding. Subsequently, 3D-ConvNets [65, 4, 15, 72, 66, 67, 16] gained popularity due to their ability to extract spatial-temporal features through 3D convolution operations. Benefiting from large-scale video datasets such as Kinetics [28], 3D-ConvNets, despite having a larger number of parameters to train, can be trained end-to-end on these extensive datasets and transferred to numerous downstream tasks. Among the commonly used 3D-ConvNets, the I3D [4] is notable for its use of RGB frames and optical flow as dual inputs to improve the extraction of motion features. On the other hand, the SlowFast Networks [16] achieve state-of-the-art results within the realm of 3D-ConvNets across many benchmark datasets. These networks require only RGB frames as input, with the Fast pathway extracting motion features that are then fused with the Slow pathway, thus saving considerable time by eliminating the need for optical flow calculations.
Recently, Transformer-based architectures such as Vision Transformer (ViT) [8], Swin Transformer [37], MLP-Mixer [64], and gMLP [34] have been known to achieve the state-of-the-art results on many visual tasks benefited with the attention layers and MLP layers which are also data-hungry. For AR, those architectures originally designed for handling images are adapted for handling the spatial-temporal dependencies in video data, including hierarchical architecture such as ViViT [1] and Video Transformer Networks (VTN) [43], and the integration of spatial and temporal attention modules such as TimeSformer [3], MViTv1 [14], and MViTv2 [31]. Finally, the Visual-Language Model (VLM) has brought attention to the research field where people are trying to train deep models for more general question-answering purposes, including querying the video data, such as Video-LLaVA [33].
@@
Action Quality Assessment. Action quality assessment (AQA) is a regression task predicting action quality scores. Most datasets, such as AQA-7 [47], Olympic Scoring Dataset [48, 50], JIGSAW [19], and MTL-AQA [49], contain sports scenes, where annotations represent the scores given by expert human judges. The typical architecture for AQA consists of 3D convolutional neural networks (C3D) [65] to extract spatial-temporal features, followed by average pooling (C3D-AVG), LSTM [25] (C3D-LSTM), or attention modules to further aggregate information from various video segments and generate regression or multitask predictions [49]. Some methods tackle this task using Label Distribution Learning [61] in which the distribution of assessment scores is predicted. Methods often predict the scores directly using MLPs [76] or the mean/std. dev of a Gaussian Distribution based on Variational Auto-Encoders (VAE) [61, 79].
@@
Skill Determination. Similar to AQA, Skill Determination [46, 45, 9, 10, 24, 32] aims to predict the skill level of the action being performed but does so via ranking. [46, 45] mainly focus on ranking surgery skills using Siamese LSTM-based networks by aggregating kinematic data. [9] investigated ranking skills for general daily activities and introduced the EPIC-Skill2018 dataset. [32] proposes an RNN-based spatial attention model and significantly improves the results in [9]. [10] employs a combination of contrastive learning and attention to discriminate between high- and low-skill actions that require fine-grained discrimination skills. Most recently, [24] proposes a TikTok dance video dataset for ranking dancing performances. Ego-Exo4D [23] is a large-scale egocentric dataset that has been released concurrent to this research, with a proficiency estimation benchmark to infer the skill level of participants from six scenarios.
@@
Video Anomaly Action Detection. Video Anomaly Detection aims to detect unintentional [11, 75] or anomalous actions [80, 59, 63, 81]. Whilst fine-grained discrimination may share some similarities to Struggle Detection, we differ in two major aspects: persons struggling may not contain anomalies, and Anomaly Detection is constructed as a temporal localisation task.
Error/Mistakes Detection. Several datasets in error/mistake detection are also conceptually related to and complementary to struggle. For example, Assembly101 [54] is known as a large-scale dataset for procedure video understanding and is also one of the first datasets to include the novel task of detecting mistakes in procedural activities. In this dataset, participants are asked to assemble LEGO toy cars while being recorded by eight static cameras capturing the process from surrounding third-person viewpoints and four egocentric cameras capturing from first-person viewpoints. The mistake actions primarily involve procedural mistakes, such as putting on a part too early and making it impossible to assemble subsequent parts, which usually require prior knowledge to determine the correct sequence. There are also skill levels of the participants annotated, ranging from 1 (worst) to 5 (best). Besides, Anomalous Toy Assembly [20] contains toy assembly videos recorded from four different viewpoints and is annotated for when the participants display sequential anomalies in the assembly task, which mainly focuses on the procedural error detection. Holoassist [69] contains multi-modality video recordings in instructor-performer pairs for training the interactive AI assistant systems. During the task performance, the instructor may capture the mistakes of the performer and intervene in the task completion procedure, including helping the performer correct the mistakes. These datasets for error/mistake detection mainly focus on procedural mistakes, which are mostly related to understanding the right sequence of steps to perform a task and usually require domain prior knowledge to be annotated. Most of the methods developed to detect procedural mistakes are similar to anomaly detection pipelines. For example, [17] proposed PREGO, in which the model is trained to anticipate the next steps from normal videos without mistakes and detect mistakes by calculating the distance from the feature of the actual videos during the test stage to see how much difference there are. This PREGO model is recognized as the first online one-class classification model for mistake detection in procedural egocentric videos and comes with two datasets adapted from Assembly101 [54] and EPIC-Tent [26]. In contrast, our struggle determination, as introduced in Section 1, though correlated with some motor errors, is more general and can be determined without prior knowledge, and in many cases, is a complement to the mistakes/errors.
@@
We find that fine-grained video understanding has been explored in a variety of different datasets and tasks (see Table 8 in the Appendix). However, despite being crucial for fine-grained activity and skill understanding, determining the participant’s level of struggle has not been attempted. Previous research efforts motivate us to explore three modelling approaches for our dataset: struggle classification, struggle level regression, and struggle label distribution learning. As labels for these tasks do not exist, we annotate struggle for one existing dataset [26, 58] and collect and annotate two new activities. Our three activities were chosen to require complex visuomotor coordination and problem-solving while remaining goal-driven. This promotes similar yet varied occurrences of struggle behaviours and provides enough diversity within and across activities.
@@ 3 Struggle Determination Dataset
We initially recorded the participants’ behaviour in three scenarios that require complex visuomotor coordination and problem-solving and in which participants may occasionally struggle to complete the tasks. These scenarios were intended to be similar to what one might encounter in daily life (both indoors and outdoors) and where one might want to receive supporting advice. Finding suitable scenarios for struggle detection presents a balancing act. The scenarios shouldn’t be too easy, as participants might not exhibit struggle behaviours, but they also shouldn’t be overly complex or open-ended to prevent unclear goals and excessive task completion times (we favour activities contained within a few minutes).
To effectively study struggle, an ideal video dataset should meet the following criteria: (1) The videos should be recorded with unobstructed views (e.g., egocentric) of a person performing a skill; (2) The dataset should contain videos of people performing a skill repeatedly; (3) The skills depicted in the video recordings should be such that they are likely to show instances of struggle.
We reviewed several existing datasets and found that most do not fully meet these criteria, particularly the second and third. Table 8 in the Appendix compares these datasets in detail.
For example, EPIC-Kitchens [7] consists of egocentric videos, satisfying the first criterion. However, the dataset primarily features daily cooking activities, which are not consistently repeated across participants, making it less suitable for studying struggle. Similarly, Ego4D Goal-Step [56] mainly captures cooking tasks, but the videos vary by recipe, making it difficult to find repeated instances of the same task with both struggling and non-struggling cases.
Some prior datasets do contain repeated activities that could be suitable for struggle annotation. For example, [26, 58] features multiple subjects repeatedly assembling a camping tent—a complex visuomotor task where struggle is more evident. We annotate this data in our work.
Other datasets, such as MECCANO [52, 51] and Assembly101 [54], contain videos of participants assembling LEGO toys. Assembly101 also includes annotations for procedural mistakes and skill levels. However, neither dataset explicitly annotates struggle, and their focus on LEGO assembly makes them narrower in scope for our study.
@@
Therefore, although many existing datasets exist, we decided to collect our own because our research focuses specifically on user struggle, which is not well-represented in those datasets. More importantly, we aimed for scenarios that, while diverse, would be ordinary enough for naive observers to estimate whether a participant was struggling without prior knowledge. To achieve this, we recorded first-person videos of two new activities, assembling plumbing pipes and game of Hanoi tower, and collected the struggle annotations of these videos from human observers. Additionally, we included the existing EPIC-Tent Dataset [26, 58], which also captures a relevant struggle-prone task of pitching a camping tent. Our dataset includes three scenarios: assembling plastic plumbing pipes by following one of two diagrams (easy or difficult) while seated indoors (Pipes), solving a version of the Tower of Hanoi puzzle while seated indoors (Tower), and assembling a camping tent while moving freely outdoors (Tent). The corresponding subsets for struggle determination are named Pipes-Struggle, Tower-Struggle, and Tent-Struggle. Each scenario consists of fine-grained actions that people perform routinely (e.g., arranging, picking up, and placing objects, attaching them, etc.). However, they differ in required equipment, difficulty, and goals. Participant performance also varies, with some struggling more than others when completing a scenario.
@@
Slightly non-struggling The actions are generally executed smoothly and with confidence, with only occasional minor indications for some level of challenge or difficulty.
Slightly struggling There are some signs of challenge or difficulty, including occasional errors, pauses, dropping off items, or wobbly hands that indicate the participants may not be very experienced in the task at hand but are still making an effort to complete it.
@@ 3.1.1 Data Collection
For all three scenarios, university undergraduate and graduate student populations were recruited. All participants gave informed consent and approval for data sharing. We recruited 24 individuals to perform the plumbing task (80% female, mean age 22.6 years), and 20 participants from this group also performed the Tower of Hanoi task. For Struggle-Tent, as stated in [26], 29 individuals (71% female, mean age 23.3 years) performed the tent-pitching scenario. In all activities, participants wore a head-mounted GoPro Hero 5 (1080p at 60Hz). Videos were captured at $1920\times 1080$ resolution, at 30Hz for Tent-Struggle, and 60Hz for Pipes-Struggle and Tower-Struggle.
@@
In the plumbing task, participants assembled home plumbing pipes following a diagram of a double sink layout (see supplementary material). Participants were evenly divided between those given an easier diagram and those given a more difficult one. To simplify dexterity requirements, the activity was reduced to assembling the layout of the pipes on a table rather than in a real sink. For Struggle-Tent, we focused on video segments from five actions in EPIC-Tent [26]: assemble support, insert stake, insert support, insert support tab, and place guyline. Our expert annotator identified these actions as having the most obvious visual cues of struggle based on a series of pilot studies.
The video data was segmented into 10-second clips, and the struggle determination annotation was accomplished by one expert and MTurkers, both of whom rated these short video clips from the three activities. We choose to annotate struggle using video-level weak labels based on the 10-second windows for the following reasons: (1) Video-level weak labels simplify the annotation process, making it easier for annotators and easier to scale. During early annotation experiments, we observed that the annotators expressed difficulty locating the beginning and end of struggle segments accurately. (2) Annotating video-level weak labels has been found to be much more time-efficient than annotating the start and end frames in long, untrimmed videos (e.g. six times more time efficient [40]. When annotators mark the start and end frames of temporal segments, they often need to replay videos multiple times to verify boundaries. In contrast, annotating small video segments with a single segment-level label for struggle reduces the need for repeated viewings. (3) Video-level weak labels help mitigate issues with the disagreement of the temporal action boundaries among the annotators, especially for more conceptual annotations, such as struggle. This was found in previous work [42], in which researchers uncovered inconsistencies in the ground truth temporal boundaries across annotators.By focusing on whether a segment contains struggle, ambiguity is reduced to determining the presence of struggle within the segment. (4) We choose 10 seconds for the size of the annotation window following evidence from Experimental Psychology research that shows that manipulation struggle actions can be identified well within 10 seconds [70, 71].
Based on the arguments above, our annotation clips thus were created by uniformly splitting the videos into 10-second segments from beginning to end (e.g. 0-9s, 10-19s, etc.). We also resized videos to $456\times 256$ at 30Hz. This ensured segments had a sufficient resolution for both human observers and the deep learning models. Our annotation process thus focuses on evaluating struggle within isolated 10-second clips. This means that we assess and label the presence of struggle solely within each clip without considering the broader context of the full video recording.